Methane at 600 psia, 100°F, is flowing in a pipe line at the rate of 10,000 cfm. Determine the mass flow rate using (a) ideal gas theory, (b) compressibility factor, (c) Redlich-Kwong equation.

1
Expert's answer
2021-12-13T16:56:14-0500

"\rho=\frac{pM}{RT}=24~\frac{kg}{m^3},"

"\dot m=\rho \dot V=113~\frac{kg}s."